PROPAGATION OF POLARIZATION IN ELASTODYNAMICS WITH RESIDUAL STRESS AND TRAVEL TIMES

SONKE HANSEN GUNTHER UHLMANN

Abstract. We show that knowing the Dirichlet-to-Neumann map (DN) associated to the equations of elastodynamics with residual stress we can determine the lens relations of presssure and shear waves. We derive several consequences of this for the inverse problem of determining the residual stress and the Lamé parameters from the DN map.